# The Impact of Domain Knowledge and Multi-Modality on Intelligent Molecular Property Prediction: A Systematic Survey

### v1

@@ 4.2.2 Semi-supervised Learning
Semi-supervised learning effectively alleviates the scarcity of labeled molecular data in fields like MPP. By blending a small subset of labeled data with a larger pool of unlabeled data, it bridges the gap between fully supervised and unsupervised learning methods. Consistency regularization is based on the idea that applying realistic perturbations to unlabeled data should not significantly alter predictions, ensuring stability and reliability in the learning process. InfoGraph*[26] employ Mean-Teacher method[162] to maximizes the mutual information between unsupervised graph representations and the representations learned by existing supervised methods in semi-supervised scenarios. Chen et al.[163] predict chemical toxicity and trained the network by the Mean Teacher SSL algorithm, which update the weights in teacher model by applying the Exponential Moving Average. Zhang et al.[27] propose a data augmentation which constructing new adjacency matrix and randomly masking the edges, and calculate the average of all data augmentation distributions and then employ MixMatch[164] label guessing and sharpening method to minimize entropy and accurately guess labels based on the label distribution center.
@@
Proxy-label strategy, assigning temporary labels to unlabeled data, expand the training dataset when labeled data is limited. This approach enhances the model’s learning process, with the proxy labels being iteratively refined for improved accuracy and generalization. ASGN[25] combines a teacher-student framework where the teacher model uses a semi-supervised approach to learn a general representation of molecular graphs and employ an active learning strategy selects informative molecules for labeling. Yu et al.[165] have developed a semi-supervised drug embedding model that combines unsupervised learning from the chemical structures of drugs and drug-like molecules with supervised learning based on hierarchical relations from an expert-crafted drug hierarchy. This approach ensures a robust and comprehensive representation of drug properties. Ma et al.[166] employ teacher-student framework, which use several epochs as a iteration, updating teacher model by the best student model. As the cross-entropy (CE) loss function is not proved to be robust to label noise during the training, they employ generalized CE[167] loss to boost the self-training. To address data imbalance, Liu et al.[168] analyze the distribution of imbalanced annotated data and identify label ranges needing adjustment, and then use high-quality pseudo-labels create graph examples to augment under-represented areas, striving for an ideal balance in training data. Wu et al.[124] introduces an instructor model to provide the confidence ratios as the measurement of pseudo-labels’ reliability. These confidence scores then guide the target model to pay distinct attention to different data points, avoiding the over-reliance on labeled data and the negative influence of incorrect pseudo-annotations.
@@
This approach not only enhances model performance by utilizing the comprehensive information available in unlabeled data but also addresses the challenge of acquiring extensive labeled datasets, a common issue in MPP.
@@ 4.2.3 Transfer Learning
Transfer learning strategies, widely adopted in various fields to address data scarcity, focus on enhancing prediction performance for tasks with limited data.[169, 170, 171] These strategies involve transferring knowledge from a data-rich source task to improve learning in a data-scarce target task. Recently, there has been a significant increase in methods employing transfer learning, showcasing its growing importance and application across different domains.
@@ 4.2.4 Multi-task Learning
Ma et al.[3] establishing a multi-label supervised model on a combined dataset with missing labels. the input to prediction network is a data matrix with multiple property label information, which can be an original dataset collected from specialized experiments. Tan et al.[32] are putting emphasis on the stacked single-target approach to make the multi-target regression and classification by various single-target base models. Biswas et al.[31] employed a multitask training method for a single model to predict critical properties and acentric factors, while also adjusting target weights in the loss function to correct data imbalance.
These works we’ve reviewed show that multi-task learning is highly effective in MPP, as it capitalizes on the interrelation of various molecular properties. This enhances a model’s capacity to simultaneously predict multiple properties, a particularly valuable trait when dealing with the challenge of limited labeled data.
@@ 5 Evaluation and Benchmark
There are a variety of benchmarks available for MPP. Notable among these are Therapeutics Data Commons, MoleculeACE[182], ADMETlab[183], USPTO-MIT[184], and MoleculeNet[33]. Each of these benchmarks offers unique datasets and challenges, allowing for a comprehensive evaluation of different models. This section should detail how models perform across these benchmarks, using a range of metrics to assess accuracy, robustness, and the efficacy of incorporating domain knowledge and multimodal data. By comparing model performances on these diverse benchmarks, we can better understand their capabilities and limitations in the complex field of MPP.
a range of benchmarks were utilized for evaluating Molecular Property Prediction (MPP), including the comprehensive MoleculeNet[33], alongside others such as MoleculeACE[182], ADMETlab[183], and USPTO-MIT[184]. Each benchmark, with MoleculeNet as a primary example, offers unique datasets and challenges, facilitating an exhaustive assessment of various models. Our analysis focuses on model performance across these benchmarks, employing metrics that gauge accuracy, robustness, and the effectiveness of integrating domain knowledge and multimodal data. This approach allows for a nuanced understanding of the strengths and weaknesses of these models in the intricate domain of MPP.
@@
Consistent with prior studies, we adopt the area under the receiver operating characteristic curve (ROC-AUC) as the evaluation metric for classification datasets, which is a widely used metric for assessing the performance of binary classification tasks. For the regression datasets, we utilize root-mean-squared error (RMSE) as the evaluation metric. It’s important to note that many studies in this field adopt either random or scaffold splits for dividing their datasets, though not uniformly. A random split involves randomly dividing the dataset into training, validation, and test sets, regardless of molecular structures. On the other hand, a scaffold split organizes molecules based on their core chemical scaffolds, ensuring that the model is tested on chemically distinct molecules from those it was trained on, providing a more stringent test of its generalization ability. The choice between these splitting methods can significantly affect the outcomes and interpretations of model performance evaluations.
@@ 6.1 Domain Knowledge Integration
Model Splitting Classification (ROC-AUC % higher is better ↑) avg. imp.
@@
As more and more work utilized the atom and bond property, the effiency of MPP is improved, but whether more atom and bond property integrate in model input? Wojtuch et al.[185] analyzed the impact of atomic features in graph convolutional neural networks, comparing twelve hand-crafted and four literature-based feature combinations. Findings indicate that feature importance is task-specific and linked to their prevalence in the dataset. Reducing less frequent or redundant features, such as formal charges or aromaticity, improves performance. These insights also apply to advanced models like graph transformers, though optimal feature selection varies by model.
Increasingly, molecular structure information is being incorporated into MPP, with several studies leveraging it to derive coarse-grained molecular insights from hierarchical graphs. Recent method like MoLGNN[18], HiGNN[77], MISU[154], CAFE[156], and iMolCLR[118] have used molecular substructure knowledge, such as BRICS or functional groups, to construct hierarchical graphs treating fragments as nodes. These methods have shown improved results over those not using substructure information. As Table 1 and Table 2 demonstrate, the ablation studies reveal a notable enhancement in methodology efficacy when fragment or functional group information is integrated. Specifically, there is a 3.98% improvement observed in regression tasks and a 1.72% improvement in classification tasks, confirming the significant impact of incorporating substructure domain knowledge in these deep learning models.
@@ 6.2 Multi-modal Data Utilization
Table 3: Comparison of multi-modal learning methods in MoleculeNet. The ‘Methods’ column specifies the learning strategy: ‘T’ denotes the use of textual data, ‘2d’ and ‘3d’ refer to the use of 2D and 3D molecular graphs, ‘CL’ indicates contrastive learning, ‘Add’ represents the direct addition fusion method, and ‘CA’ stands for cross-attention fusion. The columns—textual data, graph2d, and graph3d—highlight the types of input data utilized by the models, with the possibility of multiple selections. The best results are emboldened, and the second-best results are highlighted in red.
Model Methods T 2d 3d BBBP Tox21 ToxCast SIDER ClinTox BACE Average
KV-PLM [186] \ 72.0 70.0 55.0 59.8 89.2 78.5 70.8
GIN [187] \ 65.4 74.9 61.6 58.0 58.8 72.6 65.2
GraphMVP [18] (GIN, SchNet) CL(2d, 3d), Add 68.5 74.5 62.7 62.3 79.0 76.8 71.7
72.4 74.4 63.1 63.9 77.5 81.2 72.1
MoMu-S [188] CL(S, 2d) 70.5 75.6 63.4 60.5 79.9 76.7 71.1
MoMu-K [188] 70.1 75.6 63.0 60.4 77.4 77.1 70.6
MoleculeSTM [70] (MegaMolBART, GIN) 70.8 75.7 65.2 63.7 86.6 82.0 74.0
70.0 76.9 65.1 61.0 92.5 80.8 74.4
MolFM [101] (KV-PLM, GIN) CL(S, 2d), CA 72.2 76.6 64.2 63.2 78.6 82.6 72.9
72.9 77.2 64.4 64.2 79.7 83.9 73.7
GIT-Mol [103] (SciBERT, MoMu-S) 71.9 73.9 62.1 60.1 83.5 68.4 70.0
71.1 75.4 65.3 58.2 78.9 65.8 69.1
73.9 75.9 66.8 63.4 88.3 81.1 74.9
MolLM [189] CL(S, 2d), CL(S, 3d) 75.7 80.0 68.2 71.0 91.1 84.1 78.4
@@
This part aims to analyze the contribution of different modalities in multi-modal models for MPP. It delves into understanding how individual and combined modalities affect prediction performance. Data from various studies are collated and analyzed, emphasizing the contribution of different modalities to MPP tasks, with models in Table 3. Apart from ClinTox, there is uniformity in the predictive prowess displayed by all models across a spectrum of tasks. Nonetheless, ClinTox predictions are prone to biases in transformer-based models due to data distribution peculiarities, which result in polarized predictions. Graph-based models demonstrate a reduction in such bias, albeit with a compromised performance in ClinTox. From an input modality perspective, 3D graphical modalities exhibit superior performance. Excluding ClinTox, the performance of SMILES and graph2d is on par, with graph2d holding a slender advantage. In examining pre-training strategies, contrastive learning clearly demonstrates significant benefits. However, the integration of cross-attention might inadvertently reduce the impact of the singular modalities. Nonetheless, the strategic implementation of cross-attention promotes an effective fusion of SMILES and graph2d, resulting in combined vectors that outperform the individual modalities. In cases where modality fusion is employed, an analysis focused on the contributory significance of each data modality to the tasks reveals an interesting pattern: The standalone modalities of SMILES and graph2d yield an average performance of around 71.6%. In comparison, their fusion (SMILES+graph2d) escalates the average performance to about 74.6%. This increment attests to the potency of modality fusion, enhancing performance by an estimated 4.2%.
In conclusion, the 3D graphical modality, enriched with voluminous molecular data, is adjudged the top performer, while SMILES and graph2d exhibit comparable efficacies. In the clinTox task, addressing model bias presents a significant research opportunity. Modality fusion methods demonstrate advantages, suggesting substantial exploration potential in the ways modalities are fused.

### v3

@@ 4.2.2 Semi-supervised Learning
Semi-supervised learning effectively alleviates the scarcity of labeled molecular data in fields like MPP. By blending a small subset of labeled data with a larger pool of unlabeled data, it bridges the gap between fully supervised and unsupervised learning methods.
@@
Consistency regularization is based on the idea that applying realistic perturbations to unlabeled data should not significantly alter predictions, ensuring stability and reliability in the learning process. InfoGraph*[26] employ Mean-Teacher method[167] to maximizes the mutual information between unsupervised graph representations and the representations learned by existing supervised methods in semi-supervised scenarios. Chen et al.[168] predict chemical toxicity and trained the network by the Mean Teacher SSL algorithm, which update the weights in teacher model by applying the Exponential Moving Average. Zhang et al.[27] propose a data augmentation which constructing new adjacency matrix and randomly masking the edges, and calculate the average of all data augmentation distributions and then employ MixMatch[169] label guessing and sharpening method to minimize entropy and accurately guess labels based on the label distribution center.
@@
Proxy-label strategy, assigning temporary labels to unlabeled data, expand the training dataset when labeled data is limited. This approach enhances the model’s learning process, with the proxy labels being iteratively refined for improved accuracy and generalization. ASGN[25] adopts a teacher-student framework to jointly exploit information from molecular structure and molecular distribution to learn general representation, then employs the active learning strategy in terms of molecular diversities to select informative data. Yu et al.[170] have developed a semi-supervised drug embedding model that combines unsupervised learning from the chemical structures of drugs and drug-like molecules with supervised learning based on hierarchical relations from an expert-crafted drug hierarchy. This approach ensures a robust and comprehensive representation of drug properties. Ma et al.[171] employ teacher-student framework, which use several epochs as a iteration, updating teacher model by the best student model. As the cross-entropy (CE) loss function is not proved to be robust to label noise during the training, they employ generalized CE[172] loss to boost the self-training. To address data imbalance, Liu et al.[173] analyze the distribution of imbalanced annotated data and identify label ranges needing adjustment, and then use high-quality pseudo-labels create graph examples to augment under-represented areas, striving for an ideal balance in training data. Wu et al.[122] introduces an instructor model to provide the confidence ratios as the measurement of pseudo-labels’ reliability. These confidence scores then guide the target model to pay distinct attention to different data points, avoiding the over-reliance on labeled data and the negative influence of incorrect pseudo-annotations.
@@
This approach not only enhances model performance by utilizing the comprehensive information available in unlabeled data but also addresses the challenge of acquiring extensive labeled datasets, which is a common issue in MPP.
@@ 4.2.3 Transfer Learning
Transfer learning strategies, widely adopted in various fields to address data scarcity, focus on enhancing prediction performance for tasks with limited data.[174, 175, 176] These strategies involve transferring knowledge from a data-rich source task to improve molecular representation learning ability in a data-scarce target task. Recently, there has been a significant increase in methods employing transfer learning, showcasing its growing importance and application across different domains.
@@ 4.2.4 Multi-task Learning
Ma et al.[3] establishing a multi-label supervised model on a combined dataset with missing labels. the input to prediction network is a data matrix with multiple property label information, which can be an original dataset collected from specialized experiments. Tan et al.[32] constructed our multitask models by stacking a base regressor and classifier, enabling multitarget predictions through an additional training stage on the expanded molecular feature space. Biswas et al.[31] employed a multitask training method for a single model to predict critical properties and acentric factors, while also adjusting target weights in the loss function to correct data imbalance.
These works we’ve reviewed show that multi-task learning is highly effective in MPP, as it capitalizes on the interrelation of various molecular properties. This enhances a model’s capacity to simultaneously predict multiple properties, which is a particularly valuable trait when dealing with the challenge of limited labeled data.
@@ 5 Evaluation and Benchmark
In evaluating the performance of models in Molecular Property Prediction (MPP), it is crucial to consider a variety of benchmarks, each offering distinct datasets and posing unique challenges. Key benchmarks include MoleculeNet[33], ADMETlab[187], MoleculeACE[188], DrugOOD[189], MD17[190], TUDataset[191] (comprising MUTAG, PTC, NCI, PROTEINS, D&D, and ENZYMES), and PCQM4Mv2[192]; their details are shown in Table 1. MoleculeNet, our primary focus, offers a diverse collection of datasets in quantum mechanics, physical chemistry, biophysics, and physiology, crucial for multifaceted molecular property predictions. ADMETlab, is vital for assessing drug safety and efficacy, providing data on ADMET properties. MoleculeACE focuses on QSAR modeling challenges, notably activity cliffs. DrugOOD, based on ChEMBL, emphasizes out-of-distribution generalization in AI-aided drug discovery. MD17 is essential for validating models in computational chemistry with its molecular dynamics trajectories. TUDataset includes varied datasets like DD, ENZYMES, PROTEINS, and MUTAG, each presenting unique graph-based bioinformatics challenges. Lastly, PCQM4Mv2 from the Open Graph Benchmark offers large-scale quantum mechanical property prediction challenges for graph neural network models. Among these, MoleculeNet stands out due to its comprehensive coverage and wide usage, making it an exemplary benchmark for our evaluation.
@@
Consistent with prior studies, we adopt the area under the receiver operating characteristic curve (ROC-AUC) as the evaluation metric for classification datasets, which is a widely used metric for assessing the performance of binary classification tasks. For the regression datasets, we utilize root-mean-squared error (RMSE) as the evaluation metric.
Table 1: Overview of Datasets for Molecular Property Prediction. This table encapsulates key benchmarks, highlighting their scale, scope, and specific applications in the fields of molecular modeling, drug discovery, and computational chemistry. "Num. of Mol." means the number of molecules in the corresponding Benchmark.
Benchmark Name Description Num. of Mol. Application/Challenge
MoleculeNet[33] A diverse collection of datasets across quantum mechanics, physical chemistry, and biophysical properties, pivotal for various molecular property predictions. 785,951 Multifaceted challenges in molecular property predictions
ADMETlab[187] Provides extensive data on ADMET properties crucial for drug safety and efficacy assessments, enhancing drug development processes. 94,387 Drug development and safety evaluation
MoleculeACE[188] Focused on QSAR modeling challenges, especially activity cliffs where minor structural changes cause significant bioactivity variations, testing the robustness of ML models. 48,707 Model accuracy in subtle molecular variations
DrugOOD[189] Based on ChEMBL, it emphasizes out-of-distribution (OOD) generalization, crucial for advancing AI in drug discovery under limited and varied data scenarios. 930,314 OOD generalization in AI-aided drug discovery
MD17[190] Contains molecular dynamics trajectories, essential for developing and validating models in computational chemistry and molecular simulations. 3,817,604 Molecular dynamics model development and validation
TUDataset[191] (MUTAG, PTC, NCI, PROTEINS, D&D, ENZYMES) Includes datasets like DD, ENZYMES, PROTEINS, and MUTAG, each offering unique bioinformatics challenges in graph-based analysis, such as protein structure and enzyme function classification. —- Bioinformatics applications in graph-based learning
PCQM4Mv2[192] A dataset from the Open Graph Benchmark, providing large-scale quantum mechanical property prediction challenges for graph neural network models. 3,746,619 Quantum mechanical property prediction in molecular systems
It’s important to note that many studies in this field adopt either random or scaffold splits for dividing their datasets, though not uniformly. A random split involves randomly dividing the dataset into training, validation, and test sets, regardless of molecular structures. On the other hand, a scaffold split organizes molecules based on their core chemical scaffolds, ensuring that the model is tested on chemically distinct molecules from those it was trained on, providing a more stringent test of its generalization ability. The choice between these splitting methods can significantly affect the outcomes and interpretations of model performance evaluations.
@@ 6.1 Domain Knowledge Integration
Model Splitting Classification (ROC-AUC (%) higher is better ↑) avg. imp.
@@
As more research utilizes atom and bond properties, the efficiency of MPP has improved. However, it raises the question: does integrating additional atom and bond properties into the model input necessarily lead to higher model performance? Wojtuch et al.[193] analyzed the impact of atomic features in graph convolutional neural networks, comparing twelve hand-crafted and four literature-based feature combinations. Findings indicate that feature importance is task-specific and linked to their prevalence in the dataset. Reducing less frequent or redundant features, such as formal charges or aromaticity, improves performance. These insights also apply to advanced models like Graph Transformers, though optimal feature selection varies by model.
Increasingly, molecular structure information is being incorporated into MPP, with several studies leveraging it to derive coarse-grained molecular insights from hierarchical graphs. Recent method like MoLGNN[156], HiGNN[77], MISU[158], CAFE[160], and iMolCLR[116] have used molecular substructure knowledge, such as BRICS or functional groups, to construct hierarchical graphs treating fragments as nodes. These methods have shown improved results over those not using substructure information. As Table 2 and Table 3 demonstrate, the ablation studies reveal a notable enhancement in methodology efficacy when fragment or functional group information is integrated. Specifically, we observe a 3.98% improvement in regression tasks, measured using RMSE, and a 1.72% improvement in classification tasks, measured using ROC-AUC. These results confirm the significant impact of incorporating substructure domain knowledge into these deep learning models. We present two compelling case studies that illustrate the impact of molecular substructure information obtained via the BRICS methodology. The first example involves the molecule ’CC(C)(C)NCC(O)c1ccccc1F’. When employing BRICS fragmentation, the model identifies the fluorine atom and the tertiary amine as critical features. These fragments are known to significantly affect CNS activity due to their lipophilicity, which is a crucial determinant for blood-brain barrier (BBB) penetration. The second case focuses on CC(C)(O)C(C)(O)c1ccc(Cl)cc1, where BRICS fragmentation reveals the delicate balance between hydrophilic hydroxyl groups and lipophilic chlorinated benzene components. This balance plays a pivotal role in the molecule’s ability to penetrate the BBB. Both examples, depicted in Figure 8, showcase enhanced performance of HiGNN[77] when integrating substructure information, confirming the model’s superior ability to predict BBB penetration by capturing intricate substructure information.
@@
However, the integration of domain knowledge into molecular property prediction models is not without challenges. Firstly, there is still a lot of domain knowledge that is not digitized or gathered, even with the advances in tools like RDKit. It is possible to overlook important subtleties when converting intricate, frequently implicit expert information into an electronic format that is easy to use. Secondly, this integration process can introduce biases due to subjective interpretations by domain experts, potentially skewing model outcomes and impacting scalability and adaptability to new molecular data types. Lastly, the requirement to customize deep learning architectures to incorporate such knowledge significantly increases complexity and computational costs, complicating model development and training.
In conclusion, while domain knowledge integration is beneficial, it necessitates a careful and balanced approach. It is crucial to maintain the flexibility, scalability, and objectivity of models. These challenges highlight the need for ongoing efforts to capture and digitize comprehensive domain knowledge, maintaining a critical balance between accuracy and the practical application of these predictive models.
@@ 6.2 Multi-modal Data Utilization
Figure 8: Case study of molecular fragment information enhance for the BBBP Task. This figure compares the predictions of HiGNN[77] for two molecules, Flerobuterol and Phenaglycodol, with and without the use of BRICS-derived molecular substructure information. On the left, Flerobuterol’s molecular structure without substructure information leads to an incorrect BBBP prediction, while on the right, incorporating key fragments like the fluorine atom and tertiary amine yields an accurate prediction, highlighting these features’ role in CNS activity and BBB penetration. Similarly, for Phenaglycodol, BRICS fragmentation reveals critical hydrophilic and lipophilic components, resulting in a correct BBBP prediction, demonstrating the model’s improved predictive capability when domain knowledge is applied.
Figure 9: Case studies of multi-modal fusion for the BBBP Task. Some molecules, such as pentazocine and triamcinolone, may give incorrect predictions when based solely on 2D graph data. However, integrating SMILES information can correct these results. On the other hand, predictions based on 2D graphs for molecules like fluphenazine are accurate, whereas SMILES predictions are not. Nevertheless, combining both modalities does not compromise the overall accuracy of the final assessment. This underscores the integration of 2D graph modality data with SMILES information, which can enhance the model’s ability to correct erroneous predictions and safeguard against potential interferences.
@@
Table 4: Comparison of multi-modal learning methods in MoleculeNet. This table contrasts various models, focusing on classification (ROC-AUC %) tasks. The ‘Methods’ column specifies the learning strategy: ‘T’ denotes the use of textual data, ‘S’ denotes the use of SMILES data, ‘2d’ and ‘3d’ refer to the use of 2D and 3D molecular graphs, ‘CL’ indicates contrastive learning, and ‘CA’ stands for cross-attention fusion. The columns—SMILES data, graph2d, and graph3d—with the ✓highlight the types of input data utilized by the models, with the possibility of multiple selections. The best results are emboldened, and the second-best results are highlighted in red.
Model Methods Index Type of input data Classification (ROC-AUC % higher is better ↑) Average
S 2d 3d BBBP Tox21 ToxCast SIDER ClinTox BACE
KV-PLM [194] \ 0 72.0 70.0 55.0 59.8 89.2 78.5 70.8
GIN [195] \ 1 65.4 74.9 61.6 58.0 58.8 72.6 65.2
GraphMVP [18] (GIN, SchNet) CL(2d, 3d) 2 68.5 74.5 62.7 62.3 79.0 76.8 71.7
CL(2d, 3d), CL(2d) 3 72.4 74.4 63.1 63.9 77.5 81.2 72.1
MoMu-S [196] CL(S, 2d) 4 70.5 75.6 63.4 60.5 79.9 76.7 71.1
MoMu-K [196] 5 70.1 75.6 63.0 60.4 77.4 77.1 70.6
MoleculeSTM [58] (MegaMolBART, GIN) 6 70.8 75.7 65.2 63.7 86.6 82.0 74.0
7 70.0 76.9 65.1 61.0 92.5 80.8 74.4
MolFM [105] (KV-PLM, GIN) CL(S, 2d), CA 8 72.2 76.6 64.2 63.2 78.6 82.6 72.9
9 72.9 77.2 64.4 64.2 79.7 83.9 73.7
GIT-Mol [107] (SciBERT, MoMu-S) 10 71.9 73.9 62.1 60.1 83.5 68.4 70.0
11 71.1 75.4 65.3 58.2 78.9 65.8 69.1
12 73.9 75.9 66.8 63.4 88.3 81.1 74.9
MolLM [197] CL(T, 2d), CL(T, 3d) 13 75.7 80.0 68.2 71.0 91.1 84.1 78.4
@@
This part aims to analyze the contribution of different modalities in multi-modal models for MPP. It delves into understanding how individual and combined modalities affect prediction performance. Data from various studies are collated and analyzed, emphasizing the contribution of different modalities to MPP tasks, with models in Table 4. Apart from ClinTox, there is uniformity in the predictive prowess displayed by all models across a spectrum of tasks. Nonetheless, ClinTox predictions are prone to biases in Transformer-based models due to data distribution peculiarities, which result in polarized predictions. Graph-based models like GraphMVP, MoMu, and GIT-Mol(2d) demonstrate a reduction in such bias, albeit with compromised performance in ClinTox.
From an input modality perspective, taking the BBBP task of 2d graph and SMILES information fusion as an example from GIT-Mol[107], the size of the test dataset by scaffold split is 204. We select study cases in which the results are superior to the baseline (SciBERT) after modality fusion, as shown in Figure 9. This illustration reveals the beneficial role of SMILES modality data in augmenting graph2d data, whereby the integrated representation vectors can rectify erroneous predictions to a certain extent. Conversely, accurate predictions from graph2d, when paired with incorrect SMILES predictions, can also prevent potential mistakes, showcasing the complementary strengths of integrating diverse modalities in enhancing predictive accuracy.
In examining pre-training strategies, contrastive learning clearly demonstrates significant benefits. In examining pre-training strategies, contrastive learning clearly demonstrates significant benefits. However, the integration of cross-attention might inadvertently reduce the impact of the singular modalities. Nonetheless, the strategic implementation of cross-attention promotes an effective fusion of SMILES and graph2d, resulting in combined vectors that outperform the individual modalities.
As shown in Table 4, the methods involving modality alignment, which utilize contrastive learning between SMILES and 2D graphs, improve the model’s performance from 65.2% (index [1]) to 72.0% (average of indexes [4, 5, 7]). Furthermore, the methods integrating cross-attention mechanisms for modality fusion further enhance the model’s performance to 74.3% (average of indexes [9, 12]). GraphMVP, using contrastive learning between 2D and 3D graphs, elevates the performance from 65.2% to 71.7% (indexes [1, 2]). The MolLM achieves the optimal performance of 78.4% (index [13]) through contrastive learning and the fusion of 2D and 3D graphs.
This reveals that multi-modal learning based on 2D graphs offers a performance increase of 6.5% (71.7% - 65.2%) to 6.8% (72.0% - 65.2%) over single-modality learning, with the attention fusion mechanism providing an additional 2.3% (74.3% - 72.0%) to 6.7% (78.4% - 71.7%) boost.
In the realm of molecular property prediction, the application of multi-modality methods introduces significant challenges and limitations. Key among these is the substantial increase in computational resource consumption required for processing complex multi-modal data. This issue is particularly evident when generating detailed 3D representations from standard molecular formats like SMILES, which demands extensive resources. Concurrently, these methods often contend with processing redundant information, as a result of overlapping content among various modalities, such as SMILES, 2D graphs, and 3D structures. This overlap leads to inefficiencies due to the repeated processing of identical molecular characteristics. Additionally, integrating different data forms into a cohesive model adds another layer of complexity, necessitating a strategic approach for effective data combination to enhance predictive accuracy.
Despite these challenges, the conclusion drawn from our findings is clear: by anchoring on 2D graphs and enriching them with 1D SMILES or 3D graph information, multi-modal learning has achieved a significant ROC-AUC uplift of 9.1% to 13.2% compared to single-modality models (results from indexes [1, 9, 12, 13] in Table 4). These results underscore the substantial advantages and vast potential of modality fusion techniques in providing more holistic and comprehensive insights into molecular structures, thus enhancing the overall predictive accuracy in molecular property prediction.